At what distance a long, straight wire carrying a current of 3 A can develop a

        magnetic field of 7x10-7 T



1
Expert's answer
2022-01-23T16:14:16-0500

Answer

Let this distance is d.

Current through wire I=3A

Magnetic field B="7\times10^{-7}T"

Then distance is given by

"d=\frac{\mu_0 2 I}{4\pi B}"

Putting all values

"d=\frac{10^{-7}\times 3\times2}{7\times10^{-7}}\\=0.86m"